OpenStudy (anonymous):

Please help. I'm lost with this word Problem. (Supply and demand) A manufacturer of calculators finds that the supply and demand are given by S = 750p + 40,000 D = -400p + 132,000 (a) Find the equilibrium price. (b) What is the number of calculators sold at this price?

OpenStudy (anonymous):

S -750p = 40,000 D + 400p = 132,000 u know how to slove?

OpenStudy (anonymous):

u want me solve?

OpenStudy (anonymous):

I have no clue; i'm sorry

OpenStudy (anonymous):

frist equqtiob time (-1) = -S +750p = -40,000 -S +750p = -40,000 D + 400p = 132,000 Add------------------------- 1150p=92000 p = 92000/1150 p =80 plug p =80 find s D + 400(80) = 132,000 D+ 32000 = 132,000 D= 132000-32000 D =1000,000 here your answer

OpenStudy (anonymous):

frist equation time (-1)

OpenStudy (anonymous):

If I deserve medal then give me one

OpenStudy (anonymous):

Of course...so is the (b) at 1000,000

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

Supply =1000,000 demand =80

OpenStudy (anonymous):

Thank you so much..great..